OpenStudy (anonymous):

What is the radius of a circle defined by 2x^(2)+2y^(2)+4x-12y-30=0?

OpenStudy (anonymous):

x^2 +2x + y^2 - 6y = 15

OpenStudy (anonymous):

by diving by 3, and rearranging

OpenStudy (anonymous):

x^2 +2x + 1 +y^2 -6y + 9 = 15 + 1 + 9

OpenStudy (anonymous):

(x+1)^2 +(y-3)^2 = 25 therefore radius is sqrt25 = 5

OpenStudy (anonymous):

mean by dividing by 2

OpenStudy (anonymous):

Same type of question, just a bit more complex: Find the axes and the centre of this curve: 9x^2-4xy+6y^2=10

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (anonymous):

ahh no

OpenStudy (anonymous):

i know its from conic sections

OpenStudy (anonymous):

I think it can be done somehow with matrices

OpenStudy (anonymous):

but I am not that strong with matrix algebra

OpenStudy (anonymous):

nah i definately havent seen something like that